OpenStudy (anonymous):

A sign writer was observed painting a wall at a uniform rate. By 3pm he was 1/3 of the way up. By 4pm he was 5/6 of the way up. At what time did he reach the top?

OpenStudy (a_clan):

4:10 pm

OpenStudy (anonymous):

no

OpenStudy (anonymous):

4:20

OpenStudy (anonymous):

how

OpenStudy (anonymous):

from 3pm to 4pm he went from 1/3 to 5/6 which is equivalent to 2/6 to 5/6. That's an increase of 3/6 over 1 hour. So 1hr = 3/6 60min = 3/6 divide both sides by 3 getting 20min = 1/6 Since he only needs 1/6 more to finish, just add 20min to his last time of 4pm to get 4:20